OpenStudy (anonymous):

Prove that if the legs of a right-angle triangle are are expressible as the squares of integers, the hypotenuse cannot be an integer.

OpenStudy (anonymous):

fermat's last theorem? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

not exactly sure what the question is asking

jimthompson5910 (jim_thompson5910):

if you have something like a right triangle with side lengths of 3, 4, 5 then this is false

OpenStudy (anonymous):

but 3 is not a square of an integer

jimthompson5910 (jim_thompson5910):

oh i see

jimthompson5910 (jim_thompson5910):

Let's assume that we have a triangle with side lengths a, b, c where a = p^2 b = q^2 c = r^2 this would mean a^2 + b^2 = c^2 (p^2)^2 + (q^2)^2 = (r^2)^2 p^4 + q^4 = r^4 but yes as completeidiot is saying, there are no integral solutions to this equation by Fermat's Last Theorem

OpenStudy (anonymous):

no i was wondering whether or not it would be applicable because it doesnt state that the hypothenuse must be a square as well

jimthompson5910 (jim_thompson5910):

oh true, they just state that the hypotenuse isn't an integer at all..hmm

jimthompson5910 (jim_thompson5910):

here is one interesting fact on this page http://en.wikipedia.org/wiki/Pythagorean_triple#Elementary_properties_of_primitive_Pythagorean_triples it says that "At most one of a, b, c is a square. (See Infinite descent#Non-solvability of r2 + s4 = t4 for a proof.)"